# A Continuous Pump-Probe Experiment to Observe Rydberg Wave Packet Dynamics

### v1

@@ Abstract
Rydberg atoms remain in the limelight due to their applications in quantum optics and information technologies. In this work, the dynamics of Rydberg atoms stored in a momentum spectrometer by an all-optical trap is studied by ionizing them in the field of a continuous wave optical dipole trap. While the addition of the optical dipole trap allows to further cool the atoms, it comes at the expense of the time of flight information which is required to retrieve photoelectron momentum distributions. Here, we report on a method that extends the standard COLTRIMS (cold target recoil ion momentum spectroscopy) technique, including continuous wave lasers in pump-probe schemes, by utilizing coincidence measurements. In particular, the photoionization of atomic 6Li initially in a spin-polarized $2^{2}P_{3/2}$ state is explored. Multi-photon excitation from a tunable-mode femtosecond pulse is exploited to produce Rydberg atoms, which can then be ionized by the dipole field. The resulting ionization rate becomes explicitly time-dependent, and analyzing its structure unlocks the real-time atomic dynamics on nanosecond time scales.
@@
In this letter, we extend the toolbox for time-resolved photoelectron spectroscopy by introducing a technique that is particularly simple and suited to studying slow nanosecond dynamics. In contrast to conventional pump-probe schemes, where the time information is obtained by scanning the delay between two pulses, we use a femtosecond pulse pump and a weak continuous wave (cw) probe; the latter can photoionize the excited atomic system at any time after the former. A cold target recoil ion momentum spectrometer (COLTRIMS) Ullrich et al. (2003); Moshammer et al. (2003); Fischer et al. (2012); Fischer (2019) is employed to detect photoelectrons and recoiling target ions in coincidence. In conventional COLTRIMS, the time information of the target ionization is required to calculate the fragments’ momentum vectors. Although the atoms’ time of ionization is not directly measurable within the continuous wave, we can recover this information by exploiting the fact that the total momentum of all the target fragments is zero after photoionization (neglecting the very small photon momentum). This approach does not only retain the photoelectron energy and angular distributions, but it also provides additional information on the time evolution of the wave packet that is formed by the pump pulse. This method allows measurements of the electron’s motion as it occurs in energetically narrow wave packets, e.g., in Rydberg atoms; such systems are important in the development and study of coherent control schemes in ultra-cold atomic samples (e.g., Takei et al. (2016); Devolder et al. (2021)).
@@
For this experiment, we measure the time evolution of magnetic Rydberg wave packets in a weak external field. Atomic 6Li gas is suspended in the laser light of a near-resonant all-optical trap (AOT) Sharma et al. (2018). The AOT’s configuration optically pumps the system, resulting in a population of about 25 % in the excited $2^{2}P_{3/2}(F=5/2)$ hyperfine state. Upwards of $93\%$ of these excited atoms populate a single angular magnetic sublevel, $m_{\ell}=-1$, with respect to the quantization axis given by the spectrometer’s magnetic field and corresponding $z$-axis.
@@
To populate the Rydberg states, a femtosecond light source was used, similar to that described in Harth et al. (2017); Romans (2023), which is a commercially available optical parametric chirped-pulse amplifier (OPCPA). At its heart is a Ti:Sa oscillator which provides a broadband seed-signal for two noncollinear optical parametric amplifier (NOPA) stages. For this study, the light source was tuned to emit pulses with a central wavelength, pulse width, and repetition rate of $735\pm 10$ nm, $50$ fs, and $200$ kHz, respectively. Unlike previous experiments (e.g. De Silva et al. (2021); Acharya et al. (2021)), the intensity of this source was kept below the threshold of direct few-photon ionization with a peak intensity of up to $10^{11}$ W/cm2. This beam is intersected with the AOT at an angle of $12.5^{\circ}$ with respect to the $z$-axis. Finally, an optical dipole trap (ODT) is generated using an industrial grade ytterbium-doped fiber laser, which outputs a infrared continuous wave at $1070\pm 5$. This beam’s polarization and propagation directions lie in the $xy$-plane, and it reaches the reaction volume with an intensity of about $10^{7}$ W/cm2.
Standard reaction microscopes (ReMi) rely crucially on the timing information provided by a pulsed source. Each recoil-ion’s and electron’s time of flight $T_{r/e}$ is measured with respect to the clock set by it. Presently, the ionization is triggered by photoabsorption in a continuous field causing the time information to become lost, as atoms can be ionized at any time between pulses. To overcome this hurdle, we take advantage of aspects of the photoionization process, atomic target, and spectrometer geometry. The fragments are measured in coincidence (e.g. Moshammer et al. (2003)), which restricts their time of flight difference $\Delta T(\equiv T_{r}-T_{e})$ to a measurable, few nanosecond wide distribution. Given the low temperature of the atom cloud and the negligible momentum of the ionizing photon, we can set the total momentum of each ion-electron pair to zero Hubele et al. (2015).
@@
In detail, the charged fragments are drawn in opposite directions towards the particle detectors in the spectrometer by homogeneous electric and magnetic fields, arranged along the longitudinal $z$-axis. Thereby, the particles (charge $q$) undergo a cyclotron motion in the magnetic field $B_{z}$ due to the Lorentz force, i.e., the $x$ and $y$ positions of particles with the initial momentum components $p_{x}$ and $p_{y}$, after a time-of-flight $T$ are Fischer (2019),
@@
The key to the present time derivation method is that the fragments’ motion in the spectrometer’s electromagnetic field is highly asymmetric due to the recoil-ion being considerably more massive than its photoelectron. While their momenta are equal in magnitude, the recoil-ions’ kinetic energies are more than 5 orders of magnitude smaller than that of their electrons. It is interesting to note, that the time of flight distributions of electrons and ions (which are not directly measured in our experiment) have about the same width (around one hundred nanoseconds), while their absolute travel time can differ by more than a factor of hundred. Due to the recoil-ions’ large inertia, their cyclotron frequency $\omega_{c}(\equiv qB_{z}/m)$ is so small that the rotational phase $\omega_{c}T$, from the above equations, varies less than 1 ‰ of $2\pi$; the ions’ phase distribution can thus be approximated by a constant.
The much lighter electrons, in contrast, undergo several full cyclotron revolutions before hitting the detector, and the final rotational phases $\omega_{c}T$ depend strongly on the electrons’ initial momenta. In essence, the electrons’ rotational phase and, therefore, their time of flight is encoded in the relative positions of recoil ions and electrons on their respective detectors. If the phase variation does not exceed $2\pi$, then the electrons’ travel time can be unambiguously recovered and the three-dimensional electron momentum vectors derived.
Figure 1: Recovered photoelectron energy spectrum (top) and associated dipole-allowed energy diagram (bottom). Femtosecond and ODT photons are shown with red and purple solid arrows respectively, while the dominant spontaneous decay channels are dashed.
Details on this method can be found in Romans et al. (2023); Romans (2023), where an excellent correspondence between directly measured and derived times of flight was observed. Once recovered, this time can be used in to calculate the resultant energy and momentum distributions (see tops of Fig. 1 and Fig. 3, respectively). During an experimental run, the atoms are excited by absorption of two circulalry polarized fs-photons into a maximally aligned $f$-state wavepacket involving Rydberg states with $n\geq 8$. The excited wavepacket is then allowed to decay spontaneously back towards the ground state. During this decay, the ODT is continuously ionizing the atoms with a single photon. Energy conservation can then be used to associate peaks in the photoelectron energy distribution to internal bound states (Fig. 1, bottom).
With the time of ionization, a whole domain of time-dependent phenomena can be investigated. An example is the slow $\mu$s population dynamics driven by spontaneous decay cascades. By focusing on the set of dominant spontaneous decay channels it becomes straightforward to extract the explicit inter-state population dynamics in an effective real-time scheme Romans et al. (2023).
@@
The nanosecond accuracy also allows for the exploration of faster atomic dynamics ($\sim 1$ ns), which can be seen in Fig. 2. Here, the ionization rate is plotted as a function of photoelectron energy and the derived time of ionization. The ionization rate features a periodic time dependence throughout the whole energy range. Two dominant streaks appear, which can be associated with the $n=4$ states on the left and the Rydberg superposition on the right. Comparing with Fig. 1, the $5$D and $7$D states can be seen at around $0.6$ and $0.8$ eV, respectively; the remaining weaker lines come from the copious intermediate states in the cascade and are not shown along with the dominant decay channels.
@@
Figure 2: Derived time of ionization vs. photoelectron energy with the z-axis corresponding to ionization rate. Note how a periodic time dependence permeates all energies.
Figure 3: Semi-classical rotating magnetic moment with exaggerated tilt angle (left), initial quantum magnetic sublevel distribution (right), and the recovered photoelectron transverse momentum distribution (top). Note how the momentum ring radii correspond to the major peaks in the energy spectrum of Fig. 1.
@@
This experimental data can be explained with the simple model shown in Fig. 3. The atoms are exposed to the homogeneous magnetic spectrometer field oriented along the $z$-axis. The circularly polarized femtosecond laser beam is guided through the chamber in the $yz$-plane exciting the atoms to a Rydberg wavepacket. This momentarily maintains the alignment of the atoms’ mean magnetic moment $\expectationvalue{\vec{\mu}}$ along the photons’ propagation direction. Since this moment has a component perpendicular to the magnetic field the system will undergo Larmor precession in the $xy$-plane. Meanwhile, the atoms have sufficient energy to absorb a single infrared photon from the ODT field and be ionized. The infrared photon propagates along the $y$-axis, with its linear polarization along the $x$, shown in purple. The probability of ionizing the atom depends on the relative orientation of this photon’s polarization and the projected atomic moment. Given the geometry in the projection plane, two maxima and two minima are expected per cycle, corresponding to those times when the mean magnetic moment is aligned or perpendicular to the polarization axis, respectively.
@@
Formally, the excited wavepacket begins in aligned $f$-states, i.e. $|m_{\ell}|=3$. The alignment is tilted by $12.5^{\circ}$ with respect to the $z$-axis, i.e., the state can be expressed as a superposition of the $|m_{\textrm{z}}|\leq 3$ magnetic sublevels, as shown in Fig. 3. This excited state can be given as the following sum,
@@
with $A$ being a relative amplitude. In other words, the energy difference between the $m_{z}=-3$ and $-1$ magnetic sublevels drive the time variation of the ionization rate, with the next leading term, which depends on $m_{z}=-2$ and 00, being two orders of magnitude smaller. Since the states in the superposition have nearly identical eigenenergies, i.e. $\Omega_{m_{\textrm{z}}}^{(n)}\approx\Omega_{m_{\textrm{z}}}^{(n^{\prime})}$, they will oscillate in phase for a few microseconds.
In order to use this model and test its accuracy, the shifts of the energy state frequencies $\Omega_{m_{\textrm{z}}}^{(n)}$ must be calculated for different magnetic sublevels. Naively, one would expect that the $m_{\textrm{z}}$-depence of these shifts arises only due to the Zeeman shift, which is $m_{\textrm{z}}\cdot\omega_{L}$ in terms of the Larmor frequency $\omega_{L}(\propto B_{z})$. As it turns out, this naive assumption produces a calculated frequency two times too small, as compared to data fits.
However, besides coupling the atoms to the continuum via ionization, the ODT field can also couple lower-lying bound states via the Autler-Townes effect Autler and Townes (1955). These energy shifts can be approximated as a sum over two-state terms, $\Delta E_{AT}^{(n)}\approx\sum_{n^{\prime}}\Delta E_{n,n^{\prime}}$, for each of the dipole-allowed transitions between the $n^{\prime}D$-states and $nF$-Rydberg levels. In the large detuning limit this coupling is given as Haas et al. (2006); Hertel and Schulz (2015),
@@
Figure 4: Data and model curves for associated excited Rydberg states. The simplified single-state model is shown in red, while possible three-state superpositions are represented by the gray area.
@@
With the success of this simple model, it begs the question as to the source of the periodicity in the remaining energy levels. While the Rydberg states are excited coherently by the femtosecond laser, the lower-lying levels are populated predominately by incoherent spontaneous decay. Such a process should quickly destroy the phase information contained in the Rydberg superposition as it gets dispersed amongst the cascade. Despite this, the $n=4$ states near the $0.3$ eV peak retain a strong periodicity and seem to hold coherence throughout the experimental run.
This is explained based on relative timescales. The excited Rydberg states are comparably long-lived, and their ionization rate in the ODT field is comparably small ($\sim 0.5$ MHz); their time dependence is an expression of the states’ evolution. The $n=5,7$ states have rates roughly $10$-fold higher, while the $n=4$ states are $10$-fold higher still. This leads to the middle states having a time dependence that is a mix of the decay channels and the driving ODT field, while the short-lived states are ionized almost as soon as they are populated. Consequently, any time-dependence seen in the $n=4$ ionization data must be a delayed, but direct, projection of that seen in the Rydberg levels. The phase shift seen on the left and right sides of the $n=4$ data can also be explained by the $4P$ and $4D$ channels seen in Fig. 1; the former goes through a delayed path with the $5$ and $7$D states, while the latter has a direct dipole-allowed path to the excited Rydberg states.
@@
In conclusion, we have demonstrated a method of derived time that allows for the investigation of time-dependent atomic processes in real time. Born out of necessity, this technique utilizes the continuous field of an infrared ODT laser, in conjunction with a pulsed femtosecond pump, to constantly probe the energy states of the atoms. Lithium-6 atoms initially in a spin-polarized $2^{2}P_{3/2}$ state are excited by the fs-pump into similarly aligned $f$-states. This excitation forms a coherent superposition of Rydberg states, which then couples to the ODT’s and spectrometer’s fields. This results in shifts in the atomic energy levels via the Zeeman and Autler-Townes effects. Furthermore, the axis of alignment is slightly titled with respect to the spectrometer’s $z$-axis. This adds to the mix a superposition of magnetic sublevels or a magnetic wave packet. Each wave packet, and its associated mean magnetic moment, is dressed by the ODT’s field and undergoes an enhanced Larmor precession. In this way, the ionization rate becomes explicitly time-dependent, and through its analysis, both the relatively slow population and faster wave packet dynamics become unlocked.

### v2

# A Continuous Pump-Probe Experiment to Observe Zeeman Wave Packet Dynamics

@@ Abstract
In this work, we study the coherent dynamics of an atomic Zeeman wave packet using a continuous pump–probe scheme. A polarized wave packet is generated via few-photon excitation by a femtosecond laser pulse, creating a state with a magnetic moment tilted relative to an external magnetic field. The subsequent Larmor precession of the atoms is probed by continuous ionization in the field of an optical dipole trap (ODT) laser. Photoelectrons and photoions are detected in coincidence using a cold target recoil ion momentum spectrometer (COLTRIMS). While the addition of the ODT enables further cooling of the atomic ensemble, it removes the pulsed timing reference typically used to extract photoelectron momentum distributions in COLTRIMS. Here, we present a method that extends the standard COLTRIMS technique by exploiting redundancy in the measured data to reconstruct the time of ionization. The resulting time-dependent ionization signal reflects the coherent precession of the atomic magnetic moment, enabling real-time access to atomic dynamics on nanosecond timescales.
@@
Recently, we reported an alternative experimental approach for time-resolved photoelectron spectroscopy Romans et al. (2023), operating on the comparatively slow nanosecond timescale and thus providing a complementary perspective to the ultrafast techniques discussed above. Unlike conventional pump-probe schemes, where time information is obtained by scanning the delay between two pulses, our method employs a femtosecond pump pulse in combination with a weak continuous-wave (cw) probe. The cw probe can photoionize the excited atomic system at any time after the pump pulse, enabling time-resolved measurements without the need for delay scanning.
To detect photoelectrons and recoiling target ions in coincidence, we employ a cold target recoil ion momentum spectrometer (COLTRIMS) Ullrich et al. (2003); Moshammer et al. (2003); Fischer et al. (2012); Fischer (2019). In conventional COLTRIMS, time-of-flight information is required to reconstruct the momentum vectors of the fragments. In our approach, however, we rely solely on the measured hit positions of the particles on the detectors. A dedicated reconstruction algorithm allows us to recover not only the energy and angular distributions of the photoelectrons, but also the time of ionization relative to the femtosecond pump pulse, thereby providing additional insight into the system’s time evolution.
While our previous work Romans et al. (2023) focused on the incoherent population dynamics driven by spontaneous decay, we now investigate the coherent dynamics of the wave packet formed by the pump pulse. By using an energetically narrow laser pulse for excitation of a Zeeman wavepacket, we are able to track the electron’s motion following the excitation. Such systems are of key importance for the development and study of coherent control schemes in ultracold atomic samples (e.g., Takei et al. (2016); Devolder et al. (2021)).
@@
Specifically, we measure the time evolution of magnetic Zeeman wave packets in a weak, homogeneous external magnetic field. The general pathway is depicted in Fig. 1. First, atomic 6Li gas is suspended in the laser light of a near-resonant all-optical trap (AOT); for details of the trap, see Sharma et al. (2018). Within the AOT, the atoms are optically pumped, resulting in a population of about 25,in the excited $2^{2}P_{3/2}(F=5/2)$ hyperfine state. Of these excited atoms, more than 93,occupy a single orbital magnetic sublevel, $m_{\ell}=-1$, defined with respect to the quantization axis—the $z$-axis—which is given by the direction of the magnetic field.
@@
The excited Zeeman wave packet is generated by two-photon absorption from a femtosecond light source, which transfers population to an excited $f$-state with principal quantum number $n\geq 8$. The femtosecond laser system is similar to that described in Harth et al. (2017) and is based on a commercially available optical parametric chirped-pulse amplifier (OPCPA). At its heart is a Ti:Sa oscillator that provides a broadband seed signal for two noncollinear optical parametric amplifier (NOPA) stages. For this study, the light source was tuned to emit pulses with a central wavelength of $735\pm 10$ nm, a pulse duration of $50$ fs, and a repetition rate of $200$ kHz. Unlike in previous experiments (e.g., de Silva et al. (2021); Acharya et al. (2021)), the intensity of the source was kept below the threshold for direct few-photon ionization, with a peak intensity of up to $10^{11}$ W/cm2. The beam intersects the AOT at an angle of $12.5^{\circ}$ with respect to the $z$-axis.
Finally, the atoms are photoionized in the continuous-wave field of an optical dipole trap (ODT) laser, either directly from the highly excited state or after spontaneous decay to a lower-lying state. The ODT is generated using an industrial-grade ytterbium-doped fiber laser, which outputs an infrared continuous wave at $1070\pm 5$ nm. The laser beam’s polarization and propagation directions lie in the $xy$-plane, and it is focused into the reaction volume to a beam waist of approximately 50,$\mu$m, resulting in an intensity of about $10^{7}$ W/cm2.
Figure 1: Ionization pathway (bottom) and corresponding photoelectron energy spectrum (top). Continuous-wave AOT and ODT photons are shown in maroon and purple, respectively, while the femtosecond pulse-induced transition is depicted in red. Dominant spontaneous decay channels are indicated by dashed arrows.
The momenta of photoelectrons and photoions are measured using a cold target recoil ion momentum spectrometer. Standard COLTRIMS relies critically on time-of-flight (TOF) information for recoil ions and electrons, denoted $T_{r/e}$, which is typically measured relative to a pulsed projectile (e.g., a laser) source. In the present experiment, however, photoionization occurs in the continuous-wave field of the optical dipole trap (ODT) laser, making direct TOF measurement infeasible. Nevertheless, TOF information can be recovered by exploiting the fact that, due to the low temperature of the atomic cloud and the negligible momentum of the ionizing photon, the total momentum of each ion–electron pair can be assumed to be zero Hubele et al. (2015). As a result, the full six-dimensional momentum space of the photoelectron and photoion reduces to three independent dimensions, and their measured hit positions on the respective detectors are sufficient to reconstruct the complete kinematics of the fragmentation process.
@@
The details of the time-reconstruction method are described elsewhere Romans et al. (2023); Romans (2023); here, we provide only a brief summary. In the spectrometer, charged fragments are directed toward opposing particle detectors by homogeneous electric and magnetic fields aligned along the longitudinal $z$-axis. Due to the Lorentz force, particles with charge $q$ undergo cyclotron motion in the magnetic field, which has a single non-zero component, $B_{z}$. As a result, the $x$ and $y$ positions of particles with initial transverse momentum components $p_{x}$ and $p_{y}$ after a time-of-flight $T$ are given by Fischer (2019):
@@
Here, $\omega_{c}=qB_{z}/m$ is the cyclotron frequency of the particles. Their motion along the $z$-direction is governed by the electric field, which has its only non-zero component $E_{z}$ along this axis:
$z=\frac{p_{z}}{m}T+\frac{qE_{z}}{2m}T^{2}.$ (3)
In this equation, $z$ represents the distance between the reaction volume and the particle detector.
The key to the present time-reconstruction method lies in the fact that, although the momenta of the ion and electron are equal in magnitude (due to momentum conservation), and despite their motion being governed by the same set of equations (Eqs. (1)–(3)), their trajectories in the spectrometer’s electromagnetic field are highly asymmetric. This asymmetry arises from the significantly greater mass of the recoil ion compared to that of the photoelectron.
This becomes most apparent when considering the cyclotron phase angle from Eqs. (1) and (2)
$\omega_{c}T=\frac{B_{z}}{E_{z}\sqrt{m}}\left(\sqrt{\frac{p_{z}^{2}}{m}+2zqE_{z}}-\frac{p_{z}}{\sqrt{m}}\right),$ (4)
which is obtained by solving Eq. (3) for $T$. As can be seen from the expression on the right-hand side within the parentheses, the phase angle’s dependence on the initial momentum $p_{z}$ becomes weaker for heavier particles.
[…]
A systematic test of this method is reported in Romans et al. (2023); Romans (2023), where an excellent agreement (within 5 ns) was observed between directly measured and reconstructed times of flight. On the one hand, this timing information can be used to compute the resulting energy and momentum distributions (see the top panels of Fig.1 and Fig.3, respectively). The different photoelectron energies reflect the population distribution of the excited system following femtosecond excitation and, potentially, subsequent spontaneous decay. On the other hand, the recovered time information opens access to a broader class of time-dependent phenomena. One example is the slow, microsecond-scale population dynamics driven by spontaneous decay cascades. By focusing on the dominant decay channels, it becomes straightforward to extract explicit inter-state population dynamics in an effective real-time scheme Romans et al. (2023).
@@
The nanosecond accuracy also allows for the exploration of faster atomic dynamics ($\sim 1$ ns), which can be seen in Fig. 2. Here, the cross section of electrons being emitted in the $xy$-plane is plotted as a function of photoelectron energy and the derived time of ionization. The ionization rate features a periodic time dependence throughout the whole energy range. Two dominant streaks appear, which can be associated with the $n=4$ states on the left (centered at about 0.3 eV) and the $n\geq 8$ superposition on the right (at about 0.9 eV). Comparing with Fig. 1, the $5$D and $7$D states can be seen at around $0.6$ and $0.8$ eV, respectively; the remaining weaker lines come from the copious intermediate states in the cascade and are not shown along with the dominant decay channels.
@@
Figure 2: Ionization time delay vs. photoelectron energy, with the $z$-axis representing the rate of electron emission in the $xy$-plane.
Figure 3: Qualitative depiction of the rotating magnetic moment induced by the femtosecond excitation (left), the initial population of magnetic sublevels (right), and the corresponding recovered photoelectron transverse momentum distribution (top).
@@
This experimental data can be explained with the simple model shown in Fig. 3. The atoms are exposed to the homogeneous magnetic spectrometer field oriented along the $z$-axis. The circularly polarized femtosecond laser beam is guided through the chamber in the $yz$-plane exciting the atoms to a excited wavepacket. This momentarily maintains the alignment of the atoms’ mean magnetic moment $\expectationvalue{\vec{\mu}}$ along the photons’ propagation direction. Since this moment has a component perpendicular to the magnetic field the system will undergo Larmor precession around the $z$-axis. Meanwhile, the atoms have sufficient energy to absorb a single infrared photon from the ODT field and be ionized. The infrared photon propagates along the $y$-axis, with its linear polarization along the $x$, shown in purple. The probability of ionizing the atom and emitting photoelectrons in the $xy$-plane depends on the relative orientation of this photon’s polarization and the atom’s magnetic moment. Given the geometry, two maxima and two minima are expected per Larmor precession cycle, corresponding to those times when the mean magnetic moment is aligned or perpendicular to the polarization axis, respectively.
@@
Formally, the excited wavepacket is initially prepared in an aligned $f$-states, i.e., $|m_{\ell}|=3$ with respect to the quantization direction defined by the propagation direction of the femtosecond laser. This alignment is tilted by 12.5 relative to the laboratory $z$-axis. As a result, the wavepacket can be expressed as a coherent superposition of magnetic sublevels with $|m_{\textrm{z}}|\leq 3$, where $m_{z}$ refers to the orbital angular momentum’s projection along the $z$-axis. This decomposition is illustrated in Fig. 3. The resulting state can be written as the following sum:
@@
with $A$ being a relative amplitude. In other words, the energy difference between the $m_{z}=-3$ and $-1$ magnetic sublevels drive the time variation of the ionization rate, with the next leading term, which depends on $m_{z}=-2$ and 00, being two orders of magnitude smaller.
In order to apply this model and assess its accuracy, the energy level shifts $\Omega_{m_{\textrm{z}}}^{(n)}$ must be calculated for different magnetic sublevels. Naively, one might expect that the $m_{\textrm{z}}$-dependence of these shifts arises solely from the Zeeman effect, which contributes a term $m_{\textrm{z}}\cdot\omega_{L}$, where $\omega_{L}\propto B_{z}$ is the Larmor frequency. However, this simplistic assumption underestimates the observed frequency, and for the present magnetic field strength of approximately 4 Gauss, the calculated shift is a factor of two smaller than what is extracted from experimental data.
However, besides coupling the atoms to the continuum via ionization, the ODT field can also couple lower-lying bound states and shift the energy levels via the dynamic Stark effect. These energy shifts can be approximated as a sum over two-state terms, $\Delta E_{DS}^{(n)}\approx\sum_{n^{\prime}}\Delta E_{n,n^{\prime}}$, for each of the dipole-allowed transitions between the $n^{\prime}D$-states and $nF$-Rydberg levels. In the large detuning limit this coupling is given as Haas et al. (2006); Hertel and Schulz (2015),
@@
Figure 4: Time evolution of the ionization rate from states with principal quantum number $n\geq 8$. Experimental data points are obtained by slicing the data shown in Fig. 2, as indicated by the inset at the top. The prediction from a simplified single-state model is shown as a solid red line.
@@
With the success of this simple model, it begs the question as to the source of the periodicity in the remaining energy levels. While the states with $n\geq 8$ are excited coherently by the femtosecond laser, the lower-lying levels are populated predominately by incoherent spontaneous decay. Such a process should quickly destroy the phase information contained in the Rydberg superposition as it gets dispersed amongst the cascade. Despite this, the $n=4$ states near the $0.3$ eV peak retain a strong periodicity and seem to hold coherence throughout the experimental run.
This is explained based on relative timescales. The directly highly-excited states populated in the fs-pulse are long-lived ($\sim 1$/,$\mu$s) as compared to the lower levels ionization rates in the ODT field. The $n=5,7$ states have rates roughly $10$-fold higher, while the $n=4$ states are $10$-fold higher still. This leads to the middle states having a time dependence that is a mix of the decay channels and the driving ODT field, while the short-lived states are ionized almost as soon as they are populated. Consequently, any time-dependence seen in the $n=4$ ionization data must be a delayed, but direct, projection of that seen in the Rydberg levels. The phase shift seen on the left and right sides of the $n=4$ data can also be explained by the $4P$ and $4D$ channels seen in Fig. 1; the former goes through a delayed path with the $5$ and $7$D states, while the latter has a direct dipole-allowed path from the $n\geq 8$ states.
@@
In conclusion, we have demonstrated a method of derived time that allows for the investigation of time-dependent atomic processes in real time. This technique utilizes the continuous field of an infrared ODT laser, in conjunction with a pulsed femtosecond pump, to constantly probe the energy states of the atoms. Lithium-6 atoms initially in a spin-polarized $2^{2}P_{3/2}$ state are excited by the fs-pump into similarly aligned $f$-states. This excitation forms a coherent superposition of Rydberg states, which then couples to the ODT’s and spectrometer’s fields. This results in shifts in the atomic energy levels via the Zeeman and Autler-Townes effects. Furthermore, the axis of alignment is slightly titled with respect to the spectrometer’s $z$-axis. This adds to the mix a superposition of magnetic sublevels or a magnetic wave packet. Each wave packet, and its associated mean magnetic moment, is dressed by the ODT’s field and undergoes an enhanced Larmor precession. In this way, the ionization rate becomes explicitly time-dependent, and through its analysis, both the relatively slow population and faster wave packet dynamics become unlocked.